# ConfusionPrompt: Practical Private Inference for Online Large Language Models

### v4

@@ 4.1 Decomposer
A decomposer is designed to decompose an original user prompt into several sub-prompts, such that the required number of pseudo prompts to ensure the same level of privacy preservation can be significantly reduced. In the following, we theoretically demonstrate the benefit of a local decompose module in terms of complexity measured by the number of required pseudo prompts.
@@ 4.1.1 Complexity Analysis
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
$n\geq\left(\frac{1}{\mu}\right)^{|U(\boldsymbol{p}_{0})|}.$ (8)
@@
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompt with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
$n\geq\sum_{j=1}^{l}\left(\frac{1}{\mu}\right)^{|U(p_{0j})|}$ (9)
Specifically, if $|U(\boldsymbol{p}_{0})|=l$ and $|U(p_{0j})|=1$, $\forall j\in[1,l]$, then
@@
Theorem 11 and 12 compare the complexities between single prompt and decomposed sub-prompt, where the proof is defered to Appendix A. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. An intuitive explanation is given in Figure 2.
@@ 4.1.2 Requirements of Ideal Decomposer
  • 2.
    An ideal decomposer would optimize the complexity, i.e., the required number of pseudo prompts, in two aspects: (i) each sub-prompt contains as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
@@ 4.1.3 Model Training
We design a two-stage training approach to optimize the above objectives. In the first stage, we finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to generate decomposition given a full prompt. In the second stage, the LM is prompted to output multiple decompositions for each input. From the generations, a subset of preferred decompositions is selected to further fine-tune the model. The procedure to select preferred examples are as follows:
  • 1.
    Evaluate whether each decomposition adheres to the MECE principle, discarding any that do not meet this standard.
  • 2.
    Among the MECE-compliant decompositions, assess the complexity of each using a predefined parameter $\mu$, and select the top $k$ examples for each prompt with the lowest complexity.
@@ 4.2 Generator
Generator takes an genuine prompt $p$ and its private attributes $U(p)$ as input and produces a pseudo prompt $p^{\prime}$ with replaced attributes. The generator will be run for multiple times to obtain a series of pseudo prompts for each genuine prompt.
@@ 4.2.1 Requirements
The requirements for a pseudo prompt suggest the following criteria for an ideal generator:
  • 1.
    The generator should produce pseudo prompt with fake information that is semantically irrelevant to the correspondent attributes in genuine prompt.
  • 2.
    The pseudo prompt should not be classified as fabricated prompt by a strong discriminator $D$.
4.2.2 Model training.
The generator is trained using a similar procedure described in Section 4.1.3. In the first stage, we fine-tune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. In the second stage, we collect the preferred generations considering both semantic relevance and genuineness, with criteria given as follows:
  • 1.
    Compute $s$, the similarity between the pseudo and original prompt.
  • 2.
    Compute $f$, the genuineness score for the pseudo prompt.
  • 3.
    Calculate the overall score $\beta f-s$, where $\beta$ is the hyperparameter controlling the weight of the genuiness score.
  • 4.
@@ 4.3 Recomposer
Figure 2: Example of decomposition savings in query complexity. Decomposition module reduces the query complexity from 9 to 6 under privacy requirement $\mu=3$. This reduction becomes more pronounced with larger values of $\mu$. For instance, the complexity decreases from 100 to 20 under privacy requirement $\mu=10$.
@@ 5.1 Experiment Setup
We evaluate our framework on two multi-hop reasoning datasets: StrategyQA [36] and MuSiQue [37]. We integrate ConfusionPrompt with different online LLMs including GPT-4, GPT-4-Turbo, and GPT-4o [7]. We consider the privacy parameters $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to Appendix C.2 for the training of decomposer, generator, and recomposer.
We extract the private attributes of StrategyQA and MuSiQue through the steps below (see Appendix C.1 for details):
  • 1.
    Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. First, we extract the entities using a combination of two NER methods: Spacy [38] and Flair [39]. Then we manually correct and supplement the private attributes for each query.
  • 2.
    Step 2: model finetuning. An LLM is finetuned using the private attribute samples to guide the LLM output private attributes given a query.
  • 3.
@@ 5.2 Empirical Privacy Evaluation
We simulate two types of inference attacks to investigate the privacy protection level under various combinations of privacy parameters.
Prompt Identification Attack: an attack that identifies the true query from a group of fake and true prompts. Denote $\mathbf{q}=[\![q_{1}||q_{2}||...||q_{n}]\!]$ as the concatenated group of queries, and $k$ as the index for the true query. We finetune a BART-large classification model $G_{\rm{PIA}}$ to predict the index of the true query $\hat{k}=G_{\rm{PIA}}(\mathbf{q})$.
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [40] to predict the gender based on the user’s review.
@@ 5.3.1 Privacy Experiments
Figure 3 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$, genuineness $\rho$, similarity $\lambda$. It can be observed that: (a) decreasing significance $\mu$ consistently reduces the attack accuracy. (b) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more realistic. (c) There’s no clear relationship between the attack accuracy and similarity threshold $\lambda$. Such privacy parameter could be more related to the attribute inference attack discussed later.
Figure 3: Prompt identification attack accuracy under various combinations of privacy parameters.
Figure 4 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection metrics does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threhold $\lambda$ helps to reduce the attribute inference attack accuracies.
Figure 4: Attribute inference attack accuracy for ConfusionPrompt and LDP-based methods.
@@ 5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [41]. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against five baseline methods: (1) a local Llama2 [42] model with 7B parameters; (2) a local Vicuna [43] model with 13B parameters; (3) Text-to-text privatization (Text2Text) [30], where the query is privatized with LDP by replacing each word with the perturbed token embeddings; (4) Paraphraser [31, 32] that paraphrases the prompt with LDP mechanism using a language model; (5) querying online GPTs with no privacy protection. In the former two methods, the user downloads the open-source models and directly extract the final response from the model. In methods (3) and (4), the queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server. For the final approach, the complete prompt is transmitted to the service provider without any confusion strategy
@@
Method Privacy StrategyQA MuSiQue
Protection ACC AUC F1 roughL EM
Llama2-7B Yes 0.602 0.581 0.484 0.485 0.321
Vicuna-13B Yes 0.646 0.628 0.417 0.416 0.225
Text2Text (GPT-3.5-Turbo) Yes 0.528 0.496 0.019 0.019 0.010
Text2Text (GPT-4-Turbo) Yes 0.537 0.506 0.028 0.027 0.014
Text2Text (GPT-4o) Yes 0.533 0.500 0.021 0.020 0.012
Paraphraser (GPT-3.5-Turbo) Yes 0.489 0.478 0.076 0.076 0.028
Paraphraser (GPT-4-Turbo) Yes 0.546 0.523 0.060 0.061 0.026
Paraphraser (GPT-4o) Yes 0.537 0.515 0.072 0.072 0.038
GPT-3.5-Turbo No 0.751 0.737 0.604 0.605 0.439
GPT-4-Turbo No 0.803 0.798 0.663 0.664 0.503
GPT-4o No 0.791 0.784 0.721 0.721 0.557
ConfusionPrompt (GPT-3.5-Turbo) Yes 0.723 0.726 0.606 0.605 0.445
ConfusionPrompt (GPT-4-Turbo) Yes 0.741 0.743 0.633 0.634 0.495
ConfusionPrompt (GPT-4o) Yes 0.733 0.739 0.685 0.684 0.535
   The privacy parameters for ConfusionPrompt are fixed at $\mu=15$, $\rho=4$, and $\lambda=0.5$. The privacy budgets for LDP-based methods are set to $\epsilon=10$ and $\delta=10^{-4}$.
The results are reported in Table 1. From the results, we can see that (1) online proprietary models outperform open-source models by over 11% and 24%, repectively, for StrategyQA and MuSiQue, indicating the strong motivation for users to use online service while preserving privacy. (2) ConfusionPrompt based on various online models consistently and significantly outperforms LDP-based methods, showing the effectiveness of our proposed method. Notably, our framework possesses a distinct advantage over perturbation-based methods: increasing the privacy protection level does not adversely affect performance since the true queries are always provided to the server.
5.3.3 Decomposition Benefits
Theorem 11 and 12 show that the decomposition module could reduce the number of pseudo prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
Figure 5 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth before decomposition, whereas it exhibits approximate linearity after decomposition. Our protocol offers substantial savings ranging from an average reduction of 1.9×1.9$\times$1.9 × at $\mu=1/10$ to an average reduction of 18.3×18.3$\times$18.3 × at $\mu=1/50$.
Figure 5: Monetary ratio of strategyQA and MuSiQue dataset before (decomp) and after (w/o decomp) decompositon under various significance $\mu$. Decompostion in ConfusionPrompt substantially reduces the monetary cost, indicating its efficiency.
@@ 5.3.4 Computation Overhead Analysis
In Table 2, we present the memory cost required for the three local models, as well as that for local open-source LLMs. From the results, we can see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Specifically, it requires only $61.2\%$ and $79.7\%$ memory cost compared to running LLaMa2-7B and Vicuna 13B, respectively, on the user side.
Table 2: Memory cost for local inference with 10 samples.
Local Model Llama2-7B Vicuna-13B Decomposer Generator Recomposer Total (ConfusionPrompt)
Memory Cost 25885MB 49653MB 3867MB 3865MB 2306MB 10038MB
Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it is crucial to consider the number of sampling times it takes for a generator to produce sufficient qualified pseudo prompts. Table 3 presents the number of generations to perform and the time cost for each generation under different levels of significance $\mu$. It requires to conduct sampling for an average of $3.23$ times the $1/\mu$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute. The average time per sampling is 1.73 seconds, and thus the overall time consumption for each sample ranges from 30 seconds at $\mu=1/5$ to 161 seconds at $\mu=1/30$.
Table 3: Cost of generation in terms of number of samplings and time cost per sampling (seconds).
$1/\mu$ $5$ $10$ $20$ $30$
Number of Samplings $17.7$ $30.1$ $65.3$ $93.5$
Time per Sampling 1.73 seconds

### v5

@@ 4.1 Decomposer
A decomposer is designed to decompose an original user prompt into several sub-prompts, such that the required number of pseudo-prompts to ensure the same level of privacy preservation can be significantly reduced. In the following, we theoretically demonstrate the benefit of a local decompose module in terms of complexity measured by the number of required pseudo-prompts.
@@ 4.1.1 Complexity Analysis
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$, meaning that each contains a single query. To achieve $(\lambda,\mu,\rho)$-privacy, it requires at least $n$ prompts:
$n\geq\left(\frac{1}{\mu}\right)^{|U(\boldsymbol{p}_{0})|},$ (8)
where $|U(\boldsymbol{p}_{0})|$ represents the number of private attributes in prompt $\boldsymbol{p}_{0}$.
@@
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompts with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires at least $n$ prompts:
$n\geq\sum_{j=1}^{l}\left(\frac{1}{\mu}\right)^{|U(p_{0j})|},$ (9)
where $|U(p_{0j})|$ is the number of private attributes in the $j^{th}$ decomposed query, and $l$ denotes the number of decomposed queries in a prompt.
Specifically, if $|U(\boldsymbol{p}_{0})|=l$ and $|U(p_{0j})|=1$, $\forall j\in[1,l]$, meaning that each decomposed query contains a single attribute, then
@@
Theorem 11 and 12 compare the complexities between single prompt and decomposed sub-prompt, where the proof is defered to .1. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. An intuitive explanation is given in Figure 2, where a raw query containing three private attributes is decomposed into three sub-questions, each including a single attribute.
Figure 2: Real world example illustrating Theorem 11 and 12. The complexity is reduced from $(1/\mu)^{3}$ to $3/\mu$.
@@ 4.1.2 Requirements of Ideal Decomposer
  • An ideal decomposer would optimize the complexity, i.e., the required number of pseudo-prompts, in two aspects: (i) each sub-prompt contains as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
@@ 4.1.3 Model Training
We design a two-stage training approach to optimize the above objectives. In the first stage, we finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to generate decomposition given a full prompt. In the second stage, the LM is prompted to output multiple decompositions for each input. From the generations, a subset of preferred decompositions is selected to further finetune the model. The procedure to select preferred examples is as follows:
  • Evaluate whether each decomposition adheres to the collectively exhaustive principle, discarding any that do not meet this standard.
  • Among the qualified decompositions, assess the complexity of each using a predefined parameter $\mu$, and select the top $k$ examples for each prompt with the lowest complexity.
@@ 4.2 Generator
Generator takes an genuine prompt $p$ and its private attributes $U(p)$ as input and produces a pseudo prompt $p^{\prime}$ with replaced attributes. The generator will be run for multiple times to obtain a series of pseudo-prompts for each genuine prompt.
@@ 4.2.1 Requirements
The requirements for a pseudo-prompt suggest the following criteria for an ideal generator:
  • The generator should produce pseudo-prompt with fake information that is semantically irrelevant to the correspondent attributes in genuine prompt.
  • The pseudo-prompt should not be classified as fabricated prompt by a strong discriminator $D$.
4.2.2 Model Training
The generator is trained using a similar procedure described in Section 4.1.3. In the first stage, we finetune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. In the second stage, we collect the preferred generations considering both semantic relevance and genuineness, with criteria given as follows:
  • Compute $s$, the similarity between the fabricated and original prompt.
  • Compute $f$, the genuineness score for the pseudo-prompt.
  • Calculate the overall score $\beta f-s$, where $\beta$ is the hyperparameter controlling the weight of the genuineness score.
@@ 4.4 Algorithm
In the following, we analyze the computational complexity for the user and server.
  • User computation: For $\boldsymbol{p}_{0}=[p_{01},p_{02},...,p_{0l}]$, the computation complexity can be broken as: (i) decompose the original prompt $O(l)$; (ii) generate pseudo-prompts $O\left(1/\alpha\sum_{j=1}^{l}\left(1/\mu\right)^{|U(p_{0j})|}\right)$, where $\alpha$ denotes the probability that a generated prompt satisfies the privacy requirement. In the ideal decomposition scenario, this reduces to $O\left(|U(\boldsymbol{p}_{0})|/(\alpha\mu)\right)$. (iii) recompose the responses $O(l)$. The primary computational overhead arises from pseudo-prompt generation.
  • Server computation: The server computation complexity is $O\left(\sum_{j=1}^{l}\left(1/\mu\right)^{|U(p_{0j})|}\right)$, which, in the ideal decomposition scenario, reduces to $O\left(|U(\boldsymbol{p}_{0})|/\mu\right)$. This directly corresponds to the user query cost.
@@ 5.1 Experiment Setup
We evaluate our framework on three datasets: StrategyQA [22], MuSiQue [63], and MMLU [26]. These datasets encompass questions from a wide range of areas, such as business, medicine, and religion (see .3). We integrate ConfusionPrompt with different online LLMs including GPT-3.5-Turbo, GPT-4-Turbo, and GPT-4o [44]. We consider the privacy parameters $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to .5 for the training of decomposer, generator, and recomposer.
We extract the private attributes of the three datasets through the steps below (see .4 for details):
  • Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. Specifically, we extract the entities using a combination of two NER methods: Spacy [58] and Flair [2]. Then we manually correct and supplement the private attributes for each query.
  • Step 2: model finetuning. An LLM is finetuned using the private attribute samples so that it outputs private attributes given a query.
@@ 5.2 Empirical Privacy Evaluation
We simulate privacy attacks to investigate the privacy protection level under various combinations of privacy parameters. Existing privacy attacks can be categorized as: (i) membership inference attack that identifies whether a record is included in a model’s training dataset [57, 27], (ii) attribute inference attack that infers specific sensitive attributes [5, 4], and (iii) reconstruction attack that reconstructs the original data [74, 25].
The first attack is more associated with training time rather than inference time. Therefore, we focus on the latter two attacks, which are adapted to our ConfusionPrompt framework as follows:
Prompt Identification Attack: an attack that identifies the true query from a group of fake and true prompts. It is an adaptation of the reconstruction attack to our framework, where the attacker’s view is a set of prompts. Denote $\mathbf{q}=[\![q_{1}||q_{2}||...||q_{n}]\!]$ as the concatenated group of queries, and $k$ as the index for the true query. We finetune BART-large and Qwen3.5-4B classification models, denoted as $G_{\text{PIA}}$, to predict the index of the true query $\hat{k}=G_{\text{PIA}}(\mathbf{q})$.
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [65] to predict user gender based on the review.
@@ 5.3.1 Privacy Experiments
Figure 3 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$ and genuineness $\rho$. It can be observed that: (a) Decreasing significance $\mu$ consistently reduces the attack accuracy. (b) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more realistic. In Figure 9, we investigate the impact of similarity threshold $\lambda$ on the attack success rate, and find no clear relationship between the two (see .7). Such privacy parameter could be more related to the attribute inference attack discussed later.
Figure 3: Prompt identification attack accuracy under various combinations of significance parameter $\mu$ and genuineness parameter $\rho$ for two attack models. Similarity parameter $\lambda$ is fixed at 0.5.
Figure 4 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection metrics does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threshold $\lambda$ helps to reduce the attribute inference attack accuracies.
Figure 4: Attribute inference attack accuracy for ConfusionPrompt and LDP-based methods. Genuineness requirement $\rho$ is fixed at 4.
@@ 5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [70]. For MMLU, we report the ACC and F1 score to evaluate the multiple-choice task. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against three categories of baseline methods:
  • Local LLM. As ConfusionPrompt requires users to maintain multiple LMs for collaborative inference, it is essential to compare the performance of: (i) ConfusionPrompt utilizing online LLMs, and (ii) local inference using LLMs of comparable size to the LMs maintained on the user side. We thus evaluate the utility on two open-source LLMs: Llama2 [62] model with 7B parameters and Vicuna [11] model with 13B parameters.
  • LDP approaches. We consider two LDP approaches integrable with black-box LLMs: (i) Text-to-text privatization (Text2Text) [21], where the query is privatized with LDP by replacing each word with the perturbed token embeddings, and (ii) Paraphraser [64, 43] that paraphrases the prompt with LDP mechanism using a language model. The queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server.
  • Direct Query that queries online GPTs with no privacy protection. The complete prompt is transmitted to the service provider without any confusion strategy.
@@
Method Private StrategyQA MuSiQue MMLU
ACC AUC F1 roughL EM ACC F1
[…]
  The privacy parameters for ConfusionPrompt are fixed at $\mu=15$, $\rho=4$, and $\lambda=0.5$. The privacy budgets for LDP-based methods are set to $\epsilon=10$ and $\delta=10^{-4}$. Each value denotes the mean $\pm$ standard deviation under four rounds of experiments.
[…]
5.3.3 User Query Cost
Theorem 11 and 12 show that the decomposition module could reduce the number of pseudo-prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
Figure 5 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth in $1/\mu$ before decomposition, whereas it exhibits approximate linearity dependency on $1/\mu$ after decomposition. Our protocol offers substantial savings ranging from an average reduction of $3.3\times$ at $\mu=1/10$ to an average reduction of $18.2\times$ at $\mu=1/50$, demonstrating consistent trends in both online LLMs.
Figure 5: Monetary ratio of strategyQA, MuSiQue, and MMLU datasets for ConfusionPrompt and the framework without decomposition (ConfusionPrompt w/o Decomp) under various significance $\mu$. Decompostion in ConfusionPrompt substantially reduces the monetary cost, indicating its efficiency.
@@ 5.3.4 Computation Overhead Analysis
Memory Cost: In Figure 6, we present the memory cost required for the three local models, as well as that for local open-source LLMs. For MuSiQue, ConfusionPrompt (Base) utilizes BART-Base as the recomposer, while ConfusionPrompt (Large) employs BART-Large. For StrategyQA, ConfusionPrompt (Base) uses RoBERTa-Base as the recomposer, and ConfusionPrompt (Large) uses RoBERTa-Large.
From the results, we can see that our proposed ConfusionPrompt is significantly efficient compared to inferring large open-source language models on the user side. Even if using the large recomposer, it requires only $21.8\%$ and $41.8\%$ memory cost on average compared to run LLaMa2-7B and Vicuna 13B, respectively, on the user side. Additionally, ConfusionPrompt provides better utility than local LLM inference with lower memory requirement.
The empirical results also show the trade-off between memory requirement and utility for ConfusionPrompt. Using the base recomposer save the memory cost by an average of 17% compared with the large recomposer, with inferior performances. The suitable choice of local models depends on the user’s utility requirement and computational constraint.
Figure 6: User memory cost and utility on MuSiQue and StrategyQA for inference with 10 samples. ConfusionPrompt employs GPT-3.5-Turbo for online inference, with potential performance improvements using GPT-4-Turbo or GPT-4o.
Sampling Cost: Algorithm 1 suggests that the generator keeps generating fake prompts until the $(\lambda,\mu,\rho)$-privacy is satisfied. Therefore, it is crucial to consider the sampling time it takes for a generator to produce sufficient qualified pseudo-prompts. Figure 7 presents the number of generations to perform and the time cost for each generation under different levels of significance $\mu$. We consider three scenarios with varying combinations of genuineness parameter $\rho$ and similarity parameter $\lambda$.
For ConfusionPrompt, the sampling time scales approximately linear in significance requirement $1/\mu$. Our experiments indicate that under the most stringent privacy settings ($\rho=4$, $\lambda=0.5$), sampling is conducted an average of $2.5$ times per $1/\mu$. Relaxing the privacy constraints to $\rho=3$ and $\lambda=0.7$ reduces this ratio to approximately $2.1$. Note that $1/\mu$ can be treated as the minimal number of sampling times to achieve the privacy budget when the prompt contains only one private attribute.
Figure 7: Cost of generation in terms of time cost (seconds) for ConfusionPrompt and ConfusionPrompt w/o Decomp, under varying genuineness parameter $\rho$, similarity parameter $\lambda$, and significance parameter $\mu$.
Furthermore, we compare the generation cost for ConfusionPrompt and ConfusionPrompt w/o Decomp. We observe that decomposition reduces the sampling cost by an average of $5.88\times$ at $1/\mu=5$ and $26\times$ at $1/\mu=30$, as fewer pseudo-prompts are required to achieve the same privacy requirement.